16 October 2013

There is currently an outage in East- and West-Flanders impacting internet and telephony services. Edpnet is trying to solve this as fast as possible.

We’ll let you know as soon as any ETR is available. Sorry for the inconveniences caused.

Update 16/10/2013 12:54: Our local supplier is working on identifying the problem. We’ll keep you informed of any new information on this matter. Sorry again for the inconveniences.

Update 16/10/2013 13:05: The fault is detected: there’s a major fiber cut between Gent and Brussels. No ETR available yet, we’re cheching it with our supplier.

Update 16/10/2013 13:43: The fiber cut is finally located. We expect the splicers to arrive on location at 14:15. We’ll keep you up to date with the progress of the works and we’ll let you know the final ETR once it’s available. We apologize again for the inconvenience and thank you for your understanding.

Update 16/10/2013 14:48: The splicers are working on the issue now. Unfortunately, we still cannot provide you with a confirmed repair time. We’re monitoring the situation and will update you soon.

Update 16/10/2013 15:40: The problem seems to be quite complicated to solve, there’s almost a hundred of fiber pairs to be spliced. The ETR for the moment is around 18:00. We’ll let you know once it’s confirmed.

Update 16/10/2013 17:10: The splicing works are still ongoing. The ETR remains the same, we expect the issue to be solved at around 18:00. We’ll keep you posted.

Update 16/10/2013 18:24: The splicing works are still ongoing, got confirmation that first fibers are getting online, ours isn’t online yet. We’ll keep you posted.

Update 16/10/2013 19:42: Everything seems to be up and running again!

Thanks to everybody for your patience and understanding, we hope our next blog post will be about the new fibers being up!


19 September 2013

There is currently an outage in Ghent. It is impacting all of the internet connections in East and West Flanders. We expect the issue to be solved during the morning .

Your internet connection will re-establish itself automatically or after a reboot as soon as the problem is solved.

Update 8:40
Problem located, local Belgacom interconnect problem, no fiber outage, Belgacom ticket has been raised

Update 8:48
Belgacom acknowledges that the outage is at their side, some other clients of them are impacted. ETR for the moment is 10:00

Update 9:47
Our customers let us know they are able to surf again. Problems should now be resolved. If you can not get online, please restart your modem first.

Our apologies for the inconvenience.


8 September

There was a small interruption towards the our east- and west-flanders DSL customers, we are investigating this issue.

According to our network team, this issue should now be resolved.
We apologize for the inconvenience.


29 April 2013

There is currently an outage in East- and West-Flanders impacting internet and telephony services.
Edpnet is trying to solve this as fast as possible.

Update 29/04/2013 10:21: There’s an outage in our supplier’s network. We’ll let you know as soon as any ETR is available. Sorry for the inconveniences.

Update 29/04/2013 11:15: The fault is detected, splicers are already working on it. The connection should be restored in a few hours. We’ll keep you informed of any new information.

Update 29/04/2013 12:33: We’ve just received a confirmed ETR. The issue should be solved by 14h00.

Update 29/04/2013 13:13: It has been found out that the issue is more complicated than expected, there’s a large section of fiber to be replaced. Splicers are currently working on it. We expect it to be solved within 2 hours. Thanks a lot for your patience and understanding.

Update 29/04/2013 15:56: First of all, we’re sorry for the delay in the update. The splicing has been just finished, all customers should get online again. The end of the works is not confirmed yet though, we’ll keep you informed.

Update 29/04/2013 16:29: Our supplier confirmed that all the affected fibers were replaced. All services are up and running with no further issues since 15h55.
We apologize for the inconveniences it caused and thank you for your patience.

Edpnet apologizes for any inconveniences caused.


23 November 2012

There is currently an outage in East- and West-Flanders.
Internet and telephony is not working.
Edpnet is trying to solve this as fast as possible.

Update 24 November 2012 03:07
Fault localization continues, doesn’t seem to be any edpnet hardware issue, we try to solve this together with our supplier.  Sorry for the inconvenience.

Update 24 November 2012 08:25
Fault localization continues, a new end to end test is being performed.   Sorry for the inconvenience.

Update 24 November 2012 16:23
Fault detected at last, faulty card with supplier, a new is sent out from central stock, and should arrive within a few hours. Sorry for the inconvenience.

Update 24 November 2012 18:40
Card replaced by supplier, problem solved. Sorry for the inconvenience, and thank you for your patience.


4 October 2012

There is currently an outage in East- and West-Flanders.
Internet and telephony is not working.
Edpnet is trying to solve this as fast as possible.

Update (11:30): There is a fiber cut in the area of Gent, splicers are already on it.
Update (15:10): The fiber cut seems more complicated as first expected due to underground drilling on the location itself.  Locating the fiber cut wasn’t easy and splicers now need to dig on a difficult spot.
Meanwhile we are searching for a temporary solution for our customers, while the splicers are solving the fiber cut.  We hope to have a temporary solution by 18h00.
Update (16:55): The connection has been restored. All customers should get online again.

Edpnet apologizes for any inconveniences.


27 June 2012

Currently there is an outage in East- and West-Flanders since 13h25.
Internet and telephony is not working.
Edpnet is trying to solve this as fast as possible.

Update (14:45): The defect hardware in PoP Gent has been replaced. Everything is back online.

Edpnet apologizes for any inconveniences.


22 June 2012

A power outage occured at the main edpnet office impacting all services.

Status updates:
11:25 Start of the power outage.
11:35 Main edpnet router offline causing internet services to stop working.
11:45 Main edpnet router back online causing internet services to start working again / main edpnet office still offline.
14:00 There is currently an issue with edpnet telephony. Edpnet is trying to solve this as fast as possible.
15:00 Edpnet telephony works again as are our other services. We will closely follow-up now.
20:00 All services are up and running with no further issues since 15:00.

Edpnet apologizes for any inconveniences.


28 February 2012

Currently there is a major outage due to 2 simultaneous fiber cuts.
Internet is not working properly.
Edpnet is trying to solve this as fast as possible.

Update (13:30): 1 fiber cut has been solved, the second one is still ongoing.
Update (00:00): The second fiber cut has been solved.

Edpnet apologizes for any inconveniences.


4 January 2012

We currently have an outage at LCL Diegem.
Only colocation and IP-Transit customers are impacted.
We are trying to resolve this issue as fast as possible.

Update: The outage is solved.

Edpnet apologizes for any inconveniences.